# Breaking Barriers: Do Reinforcement Post Training Gains Transfer To Unseen Domains?

### v2

@@ 4 Findings
  • (RQ2) RPT demonstrates cross-domain generalizability when reasoning patterns are similar, such as mutual transfer between math and code, but fails to generalize across distinct reasoning patterns, such as from math or code to general reasoning (Section 4.2).
@@ 4.1 RPT Gains Do Not Generalize to Arbitrary Unseen Domains
We present the results in Table 2. Across the table, RPT models exhibit considerably higher improvements on ID tasks compared to OOD tasks, with a 3.57% increase in pass@1 for ID tasks, but a 1.48% decrease for OOD tasks. For example, (1) DeepScaleR-1.5B-Preview shows a 5.1% gain in pass@1 on math domain tasks, but only 1.7% in others, representing a 3×3$\times$3 × drop. This lack of generalizability stems from the RPT algorithm itself rather than simply from overfitting to large-scale training data: notably, a similar trend is observed in (6) Absolute_Zero_Reasoner-Coder-7B, which was fine-tunedd on a near-zero amount of data. Despite its minimal training data exposure, this model experiences a 23.31% decrease in pass@1 accuracy on unseen domains, while achieving a 30.12% improvement within its RPT domain. We also observe that the generalizability of RPT algorithms is sensitive to the training data, implementation details, and finetuning strategy. For example, although all trained on math data, (1) DeepScaleR-1.5B-Preview demonstrates improvements in both $ID$ and $OOD$ tasks, whereas (7) ZR1-1.5B and (10) STILL-3-1.5B-Preview show statistically significant performance gains in $ID$ tasks as well as statistically significant performance drops on $OOD$ tasks. These findings suggest that the gains from RPT are largely domain-specific: models significantly improve on tasks similar to their training data, but fail to generalize robustly to other unseen domains.
@@
Table 2: Existing Open-sourced RPT models achieve significantly larger accuracy gains $\Delta$ (%) and odds ratios $\hat{\theta}$ on in-domain (ID) tasks compared to out-of-domain (OOD) tasks. An asterisk (*) denotes statistical significance at $p<0.05$.
Metric (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) (14) Avg.
$\Delta^{(ID)}\uparrow$ 5.40 4.61 6.96 30.03 -6.27 30.12 2.82 7.07 -26.01 4.31 -4.27 -3.84 -0.41 -0.59 3.57
$\Delta^{(OOD)}\uparrow$ 1.67 4.01 -2.64 46.32 2.55 -23.31 -5.47 13.22 -4.73 -1.33 -7.39 -27.89 -8.44 -0.34 -1.48
$\Delta^{(ID)}-\Delta^{(OOD)}$ 3.73 0.60 33.37 -16.29 -8.82 53.43 8.30 -6.15 -2.13 5.64 3.12 24.05 -8.85 -0.25 5.04
$\hat{\theta}^{(ID)}\uparrow$ $1.36^{*}$ $1.45^{*}$ $1.59^{*}$ $7.13^{*}$ $0.52^{*}$ $22.47^{*}$ $1.22^{*}$ $1.34^{*}$ $0.34^{*}$ $1.28^{*}$ $0.72^{*}$ $0.83^{*}$ $0.97$ $0.97$ 3.01
$\hat{\theta}^{(OOD)}\uparrow$ $1.07^{*}$ $1.18^{*}$ $0.31^{*}$ $14.09^{*}$ $1.15^{*}$ $0.41^{*}$ $0.80^{*}$ $1.98^{*}$ $0.68^{*}$ $0.95^{*}$ $0.69^{*}$ $0.30^{*}$ $1.47^{*}$ $0.99$ 1.86
$\hat{\theta}^{(ID)}/\hat{\theta}^{(OOD)}$ 1.27 1.22 5.15 0.51 0.45 54.61 1.53 0.68 0.50 1.35 1.03 2.74 0.66 0.98 5.19
@@
Figure 2: RPT models on single domains show significant pass@1 improvements over base models and higher odds ratios on in-domain tasks, but not on out-of-domain tasks. No single-domain model achieves statistically significant gains in out-of-domain tasks (* denotes $p<0.05$).
@@
Single-domain finetuning reinforces evidence of RPT’s limited generalizability. To further dissect the generalizability limitations identified in our observational study, we conduct a more controlled investigation by isolating models fine-tuned exclusively on single domains. To do so, we analyze our interventional study results, where $ID$ corresponds to the training domain, while $OOD$ include all tasks from the remaining two domains in our evaluation.
@@
As shown in Figure 2, none of the models fine-tunedd on a single domain exhibit statistically significant improvement on $OOD$ tasks. Both the Math-RPT model and the Code-RPT model show performance drops on $OOD$ tasks with statistical significance, in contrast to the statistically significant gains they achieve in-domain. The General-RPT model also demonstrates no statistically significant gains on its $OOD$ tasks.
@@ 4.2 RPT Gains Generalize Across Domains with Similar Reasoning Patterns
(b) Odds ratio $\hat{\theta}^{(\mathcal{D})}$ across domains. An asterisk (*) denotes statistical significance at $p<0.05$.
Figure 3: Multi-domain evaluation results of existing RPT models. We highlight in-domain results with frames. RPT shows mutual generalizability between math and code, one-way transfer from general reasoning to math and code, but no generalization from math or code to general reasoning.
Structured-to-structured generalization is effective. We observe that models fine-tuned on math and code data exhibit strong mutual generalizability. In our observational analysis (Figure 3), models fine-tuned exclusively on math or code demonstrate transferable performance gains across these two domains. For example, models fine-tuned on math domain data achieve an average improvement of 2.18% in pass@1 on math domain tasks and 4.77% on code domain tasks. Similarly, models fine-tuned on code domain data improve by 9.49% in pass@1 on code domain tasks and 15.44% on math domain tasks. In both cases, the improvement is even greater on the non-finetuned domain, suggesting that math and code tasks share common structured reasoning patterns that enable RPT to generalize effectively across these domains.
@@
Figure 4: Multi-domain evaluation results of RPT models on single domains. An asterisk (*) denotes statistical significance at $p<0.05$. We highlight in-domain results with frames. RPT demonstrates generalizability from math to code and from general reasoning to math, but shows no generalizability from math or code to general reasoning.
@@
Structured reasoning patterns that are more foundational tend to exhibit stronger cross-domain transfer. Building on the findings from our observational study, we further examine the generalizability across structured reasoning domains using interventional study results from models fine-tuned on single domains (Figure 4). We observe that the generalizability from math to code is notably stronger and more consistent than the reverse. This aligns with the intuition that mathematical reasoning is a more fundamental form of structured thinking, serving as the backbone for coding tasks, and thus enables better cross-domain transfer when used as RPT data.
@@
Structured-to-unstructured generalization is limited. Models trained on structured reasoning domains—such as math and code—exhibit substantially reduced improvements when evaluated on general-domain tasks. In our observational study (Figure 3), models fine-tuned on structured reasoning domains (i.e., math, code, or both) achieve an average improvement of -0.27% in pass@1 on general-domain tasks, compared to significant gains of 11.08% on math and 5.82% on code tasks. While the improvements in math and code domain tasks are statistically significant, the performance drops on general reasoning tasks, indicating a lack of generalizability to unstructured domains. For instance, in the (1) DeepScaleR-1.5B-Preview and (2) DeepCoder-1.5B-Preview pair, the observed gains in math and code domain tasks are both significantly higher than those in general-domain tasks. Our interventional study results further confirm this trend (Figure 4): while the Math-RPT model shows improvements in both math and code domain tasks, its performance drops notably on general-domain tasks. Similarly, the Code-RPT model shows a statistically significant drop in performance on general reasoning tasks. These results suggest that although structured reasoning skills generalize well across similarly structured domains, they fail to transfer effectively to domains that require less structured, more heterogeneous reasoning patterns.
@@
Unstructured-to-structured generalization is promising. RPT models trained on unstructured general-domain data still exhibit measurable gains in structured tasks. In our observational study (Figure 3), general-domain RPT models show substantially higher pass@1 improvements on math (21.40%) and code (12.16%) tasks compared to tasks within the general reasoning domain. Similarly, in our interventional analysis (Figure 4), the General-RPT model achieves statistically significant gains on math domain tasks and shows no noticeable degradation on code domain tasks, while underperforming on tasks within its own domain. This suggests that unstructured reasoning patterns encompass broader representational complexity and implicitly subsume the essential components of structured reasoning, functioning as a conceptual superset.
4.3 Intra-domain RPT Gains Depend on Structural Similarity among Subdomains
@@
Structured reasoning patterns generalize effectively within domain. Consistent with prior work on math and code reasoning, our observational study shows that models fine-tunedd on structured domains generalize well across tasks within the same domain (Figure 3). On average, models trained on math data achieve a pass@1 improvement of 2.18% on math tasks, while models trained on code data show an average improvement of 9.49% on code tasks. Our interventional analysis further confirms this trend where structured-domain models (i.e., the Code-RPT model and the Math-RPT model) exhibit the largest gains on tasks from their corresponding training domain (Figure 4). These results suggest that data following consistent and structured reasoning templates facilitates reliable generalization within the same domain, as downstream tasks can leverage similar inductive patterns.
@@
Unstructured reasoning patterns lack intra-domain consistency. In contrast, models trained on general-domain (unstructured) data demonstrate limited or negative transfer to other unstructured tasks from different domains in our observational study (Figure 3). For instance, Fino1-8B (model (12)), fine-tunedd on financial data, exhibits notable performance drops when evaluated on all unrelated general-domain tasks. Its pass@1 on PubMedQA (medical domain) declines from 3.26% to 1.28%, on LegalBench (legal domain) from 6.42% to 4.84%, and on TabFact declines from 64.18% to 48.39% (general tabular knowledge). Our interventional results reinforce this observation: the General-RPT model underperforms the base model on general-domain tasks, with the degradation in accuracy being statistically significant (Figure 4). This suggests that, unlike structured domains, unstructured reasoning tasks are highly diverse and domain-specific. They lack a shared logical template, making it difficult for RPT to generalize even within what is nominally the same domain.
@@ 5 Related Work
In this section, we review RPT from the following four aspects.
RPT algorithms. The dominant approach for RPT is Proximal Policy Optimization (PPO) [56, 55], which builds on classical policy-gradient methods such as REINFORCE [72] and Trust Region Policy Optimization (TRPO) [55] and has been widely used in the post-training of OpenAI’s ChatGPT [49], Google’s Gemini [65], and Anthropic’s Claude [5]. However, PPO is an actor-critic-based algorithm, which requires an additional critic network that is typically also initialized from the same pretrained LLM. This introduces both computational overhead and algorithmic complexity compared to the vanilla REINFORCE algorithm [73]. In recognition of this, recent work has attempted to design simplified alternative approaches to PPO for LLM post-training. A line of work studies arguably simplest rejection sampling fine-tuning [17, 69], which iteratively generates multiple completions per prompt, filters out low-quality responses, and fine-tunes on the selected outputs. Another direction revisits Reinforce-style methods, such as GRPO [58], DAPO [82], VAPO [85], ReMax [35], RLOO [2, 31], Reinforce++ [24], and Reinforce-rej [76]. These algorithms discard the critic network and instead rely on Monte Carlo estimates from on-policy samples to update the policy. These methods primarily differ in how they estimate the advantage function. Among them, GRPO has garnered particular attention for its strong empirical performance in the post-training of DeepSeek-R1 [57].
RPT frameworks. RPT was first widely applied under the reinforcement learning from human feedback (RLHF) framework in the context of LLM post-training [52, 8]. In RLHF, a proxy reward is learned from human-annotated preference dataset and Bradley-Terry model [10], and the models are trained to optimize a KL-regularized reward objective to prevent overfitting the imperfect learned reward. RLHF has since become a standard technique in LLM training pipelines [49, 5, 69, 65]. More recently, RPT with verifiable reward has received significant attention in building powerful reasoning models, following the release of OpenAI o1 [51] and DeepSeek-R1 [16]. In this framework, a verifier is employed to check the correctness of the final answer for reasoning-related tasks, and serves as the reward signal for the RPT training. The major advantage is that this verifier score is much more reliable than the learned reward in RLHF, thus enabling large-scale RPT training in the post-training stage. This has also led to the development of increasingly capable and open-source RPT frameworks [60, 39, 40].
Limitations of RPT. Despite the recent successes of RPT in improving language model reasoning capabilities, the limitations of RPT in general have been widely studied. At the training phase, SFT with LLM reasoning traces, without RL, has been shown to be effective enough [67, 84]. At the inference phase, the quality of reasoning models depends crucially on their ability to scale under test-time compute constraints [47, 88]. Moreover, the effectiveness of LLMs’ lengthy "thinking" processes has also been challenged: recent findings suggest that bypassing explicit multi-step reasoning through simpler prompting or parallelized sampling can achieve comparable or even superior results [42, 81]. Building on these observations, our work aims to examine the limitations of RPT at a finer granularity, specifically its data generalizability across the training and inference phases.
Applications of RPT across tasks. RPT has proven effective for a broad range of tasks with well-defined correctness, where LLMs are finetuned using reward feedback. These range from structured tasks such as math [77, 14] and coding [71, 62, 32, 59], to unstructured tasks like search engine use [29] and open-ended question answering [63]. However, all these models are trained and evaluated on tasks within a single domain or task type. Even works on general knowledge [63] remain confined to open-ended question answering tasks, without testing transfer across fundamentally different task types. In contrast, our work directly addresses this cross-domain generalization gap.

### v3

@@ 4 Findings
  • (RQ2) RPT demonstrates cross-domain generalizability when reasoning patterns are similar, such as mutual transfer between math and code, but fails to generalize across distinct reasoning patterns, such as from math or code to knowledge-intensive reasoning (Section 4.2).
@@
  • (RQ4) The generalizability of RPT is consistent across different hyperparameters (Section 4.4).
@@ 4.1 RPT Gains Do Not Generalize to Arbitrary Unseen Domains
We present the results in Table 2. Across the table, RPT models exhibit considerably higher improvements on ID tasks compared to OOD tasks, with a 2.87% increase in pass@1 for ID tasks, but a 3.19% decrease for OOD tasks. For example, (1) DeepScaleR-1.5B-Preview shows a 5.1% gain in pass@1 on math domain tasks, but only 1.7% in others, representing a $3\times$ drop. This lack of generalizability stems from the RPT algorithm itself rather than simply from overfitting to large-scale training data: notably, a similar trend is observed in (6) Absolute_Zero_Reasoner-Coder-7B, which was post-trained on a near-zero amount of data. Despite its minimal training data exposure, this model experiences a 23.31% decrease in pass@1 accuracy on unseen domains, while achieving a 30.12% improvement within its RPT domain. We also observe that the generalizability of RPT algorithms is sensitive to the training data, implementation details, and finetuning strategy. For example, although all trained on math data, (1) DeepScaleR-1.5B-Preview demonstrates improvements in both $ID$ and $OOD$ tasks, whereas (7) ZR1-1.5B and (10) STILL-3-1.5B-Preview show statistically significant performance gains in $ID$ tasks as well as statistically significant performance drops on $OOD$ tasks. These findings suggest that the gains from RPT are largely domain-specific: models significantly improve on tasks similar to their training data, but fail to generalize robustly to other unseen domains.
@@
Table 2: Existing Open-sourced RPT models achieve significantly larger accuracy gains $\Delta$ (%) and odds ratios $\hat{\theta}$ on in-domain (ID) tasks compared to out-of-domain (OOD) tasks.
Metric (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) (14) (15) (16) (17) (18) Avg.
$\Delta^{(ID)}\uparrow$ 5.40 4.61 6.96 11.62 -6.27 30.12 2.82 7.07 -26.01 4.31 -4.27 -3.84 -0.41 -0.59 -1.87 0.61 -3.60 25.01 2.87
$\Delta^{(OOD)}\uparrow$ 1.67 4.01 -26.41 -3.70 2.55 -23.31 -5.47 13.22 -4.73 -1.33 -7.39 -27.89 8.44 -0.34 -5.68 -0.04 -5.43 24.34 -3.19
$\Delta^{(ID)}-\Delta^{(OOD)}$ 3.73 0.60 33.37 15.32 -8.82 53.43 8.30 -6.15 -2.13 5.64 3.12 24.05 -8.85 -0.25 3.81 0.65 1.83 0.67 6.07
$\hat{\theta}^{(ID)}\uparrow$ $1.36^{*}$ $1.45^{*}$ $1.59^{*}$ $2.37^{*}$ $0.52^{*}$ $22.47^{*}$ $1.22^{*}$ $1.34^{*}$ $0.34^{*}$ $1.28^{*}$ $0.72^{*}$ $0.83^{*}$ $0.97$ $0.97$ $0.87^{*}$ 1.06 $0.85^{*}$ $15.58^{*}$ 3.10
$\hat{\theta}^{(OOD)}\uparrow$ $07^{*} 1.1$$18^{*}$ $0.31^{*}$ $0.86^{*}$ $1.15^{*}$ $0.41^{*}$ $0.80^{*}$ $1.98^{*}$ $0.68^{*}$ $0.95^{*}$ $0.69^{*}$ $0.30^{*}$ $1.47^{*}$ $0.99$ $0.71^{*}$ 1.00 $0.70^{*}$ $8.50^{*}$ 1.32
$\hat{\theta}^{(ID)}/\hat{\theta}^{(OOD)}$ 1.27 1.22 5.15 2.75 0.45 54.61 1.53 0.68 0.50 1.35 1.03 2.74 0.66 0.98 1.22 1.06 1.20 1.83 4.46
@@
Figure 2: RPT models on single domains show significant pass@1 improvements over base models and higher odds ratios on in-domain tasks, but not on out-of-domain tasks. No single-domain model achieves statistically significant gains in out-of-domain tasks.
@@
Single-domain finetuning reinforces evidence of RPT’s limited generalizability. To further dissect the generalizability limitations identified in our observational study, we conduct a more controlled investigation by isolating models post-trained exclusively on single domains. To do so, we analyze our interventional study results, where $ID$ corresponds to the training domain, while $OOD$ include all tasks from the remaining two domains in our evaluation.
@@
As shown in Figure 2, none of the models post-trained on a single domain exhibit statistically significant improvement on $OOD$ tasks. Both the Math-RPT model and the Code-RPT model show performance drops on $OOD$ tasks with statistical significance, in contrast to the statistically significant gains they achieve in-domain. The Knowledge-RPT model also demonstrates no statistically significant gains on its $OOD$ tasks.
@@ 4.2 RPT Gains Generalize Across Domains with Similar Reasoning Patterns
(b) Odds ratio $\hat{\theta}^{(\mathcal{D})}$ across domains.
Figure 3: Multi-domain evaluation results of existing RPT models. We highlight in-domain results with frames. RPT shows mutual generalizability between math and code, one-way transfer from knowledge-intensive reasoning to math and code, but no generalization from math or code to knowledge-intensive reasoning.
Structured-to-structured generalization is effective. We observe that models post-trained on math and code data exhibit strong mutual generalizability. In our observational analysis (Figure 3), models post-trained exclusively on math or code demonstrate transferable performance gains across these two domains. For example, models post-trained on math domain data achieve an average improvement of 2.18% in pass@1 on math domain tasks and 4.77% on code domain tasks. Similarly, models post-trained on code domain data improve by 9.49% in pass@1 on code domain tasks and 15.44% on math domain tasks. In both cases, the improvement is even greater on the non-finetuned domain, suggesting that math and code tasks share common structured reasoning patterns that enable RPT to generalize effectively across these domains.
@@
Figure 4: Multi-domain evaluation results of single domain RPT models. We highlight RPT domains with frames. RPT demonstrates generalizability from math to code and from knowledge-intensive reasoning to math, but shows no generalizability from math or code to knowledge-intensive reasoning.
@@
Structured reasoning patterns that are more foundational tend to exhibit stronger cross-domain transfer. Building on the findings from our observational study, we further examine the generalizability across structured reasoning domains using interventional study results from models post-trained on single domains (Figure 4). We observe that the generalizability from math to code is notably stronger and more consistent than the reverse. This aligns with the intuition that mathematical reasoning is a more fundamental form of structured thinking, serving as the backbone for coding tasks, and thus enables better cross-domain transfer when used as RPT data. We provide a quantitative analysis of reasoning trace similarity across structured domains in Appendix B.1.
@@
Structured-to-unstructured generalization is limited. Models trained on structured reasoning domains, such as math and code, exhibit substantially reduced improvements when evaluated on knowledge-intensive domain tasks. In our observational study (Figure 3), models post-trained on structured reasoning domains (i.e., math, code, or both) achieve an average improvement of -0.27% in pass@1 on knowledge-intensive domain tasks, compared to significant gains of 11.08% on math and 5.82% on code tasks. While the improvements in math and code domain tasks are statistically significant, the performance drops on knowledge-intensive reasoning tasks, indicating a lack of generalizability to unstructured domains. For instance, in the (1) DeepScaleR-1.5B-Preview and (2) DeepCoder-1.5B-Preview pair, the observed gains in math and code domain tasks are both significantly higher than those in knowledge-intensive domain tasks. Our interventional study results further confirm this trend (Figure 4): while the Math-RPT model shows improvements in both math and code domain tasks, its performance drops notably on knowledge-intensive domain tasks. Similarly, the Code-RPT model shows a statistically significant drop in performance on knowledge-intensive reasoning tasks. These results suggest that although structured reasoning skills generalize well across similarly structured domains, they fail to transfer effectively to domains that require less structured, more heterogeneous reasoning patterns.
@@
Unstructured-to-structured generalization is promising. RPT models trained on unstructured knowledge-intensive domain data still exhibit measurable gains in structured tasks. In our observational study (Figure 3), knowledge-intensive domain RPT models show substantially higher pass@1 improvements on math (21.40%) and code (12.16%) tasks compared to tasks within the knowledge-intensive reasoning domain. Similarly, in our interventional analysis (Figure 4), the Knowledge-RPT model achieves statistically significant gains on math domain tasks and shows no noticeable degradation on code domain tasks, while underperforming on tasks within its own domain. This suggests that unstructured reasoning patterns encompass broader representational complexity and implicitly subsume the essential components of structured reasoning, functioning as a conceptual superset.
4.3 Intra-domain RPT Gains Depend on Subdomain Structural Similarity
@@
Structured reasoning patterns generalize effectively within domain. Consistent with prior work on math and code reasoning, our observational study shows that models post-trained on structured domains generalize well across tasks within the same domain (Figure 3). On average, models trained on math data achieve a pass@1 improvement of 2.18% on math tasks, while models trained on code data show an average improvement of 9.49% on code tasks. Our interventional analysis further confirms this trend where structured-domain models (i.e., the Code-RPT model and the Math-RPT model) exhibit the largest gains on tasks from their corresponding training domain (Figure 4). These results suggest that data following consistent and structured reasoning templates facilitates reliable generalization within the same domain, as downstream tasks can leverage similar inductive patterns.
@@
Unstructured reasoning patterns lack intra-domain consistency. In contrast, models trained on knowledge-intensive domain (unstructured) data demonstrate limited or negative transfer to other unstructured tasks from different domains in our observational study (Figure 3). For instance, Fino1-8B (model (12)), post-trained on financial data, exhibits notable performance drops when evaluated on all unrelated knowledge-intensive domain tasks. Its pass@1 on PubMedQA (medical domain) declines from 3.26% to 1.28%, on LegalBench (legal domain) from 6.42% to 4.84%, and on TabFact declines from 64.18% to 48.39% (general tabular knowledge). Our interventional results reinforce this observation: the Knowledge-RPT model underperforms the base model on knowledge-intensive domain tasks, with the degradation in accuracy being statistically significant (Figure 4). This suggests that, unlike structured domains, unstructured reasoning tasks are highly diverse and domain-specific. They lack a shared logical template, making it difficult for RPT to generalize even within what is nominally the same domain.
We provide quantitative evidence for intra-domain diversity of different domains in Appendix B.2.
4.4 RPT Generalizability Remains Weak Across Configuration Variants
RPT generalizability is consistent across different algorithms. By comparing the in-domain and out-of-domain performance differences between GRPO and DAPO on the math domain (Table 3), we observe that both algorithms yield similar gain gaps. This suggests that despite their procedural differences, the core optimization behavior of RPT dominates, leading both algorithms to learn essentially the same domain-specific reasoning patterns.
RPT generalizability is consistent across different base models. By comparing the in-domain and out-of-domain performance differences between base models of DeepSeek-R1-Distill-Qwen-1.5B and Llama-3.2-3B-Instruct on the math domain (Table 3), we observe that both base models yield similar gain gaps. This suggests that the lack of generalizability in RPT is inherent to the RPT process itself, rather than a consequence of the base model architecture or pretraining data.
Table 3: RPT configuration variants consistently fail to improve generalizability, as shown by accuracy gains $\Delta$ (%) and odds ratios $\hat{\theta}$ on in-domain (ID) tasks versus out-of-domain (OOD) tasks across different base models and RPT algorithms.
Base Model + RPT Algorithm $\Delta^{(ID)}\uparrow$ $\Delta^{(OOD)}\uparrow$ $\Delta^{(ID)}-\Delta^{(OOD)}$ $\hat{\theta}^{(ID)}\uparrow$ $\hat{\theta}^{(OOD)}\uparrow$ $\hat{\theta}^{(ID)}/\hat{\theta}^{(OOD)}$
DeepSeek-R1-Distill-Qwen-1.5B + GRPO 3.13 -1.81 4.94 $1.1881^{*}$ $0.9269^{*}$ 1.2812
Llama-3.2-3B-Instruct + GRPO 6.47 1.41 5.06 $1.4694^{*}$ $1.0635^{*}$ 1.3817
DeepSeek-R1-Distill-Qwen-1.5B + DAPO 3.96 -1.27 5.23 $1.2395^{*}$ $0.9482^{*}$ 1.3070
RPT generalizability decreases as training steps increase. As we show in Figure 5, the gap between in-domain and out-of-domain gains increases as the number of training steps increases and eventually stabilizes. This indicates that model generalizability gradually declines as RPT progresses. The reason is that increased exposure to the domain-specific data leads the model to overfit, improving in-domain performance while offering diminishing gains out-of-domain. The eventual convergence from a full epoch onwards occurs because the model is repeatedly trained on the same finite collection of data, limiting further changes in the gap betwen in-domain and out-of-domain gains. We conduct more fine-grained evaluations and analyses in Appendix H.
(a) Pass@1 improvement $\Delta^{(\mathcal{D})}$ (%).
(b) Odds ratio $\hat{\theta}^{(\mathcal{D})}$ across domains.
Figure 5: In-domain and out-of-domain improvements during RPT training on the math domain. The gap between in-domain and out-of-domain improvements grows as training progresses.
Larger model sizes do not lead to better generalizability. By comparing performance differences across model-size variants within the same model family, we find that the average in-domain gain increases by 16.5% more than the out-of-domain gain as model size grows. This pattern is consistent with the intuition that larger models are more prone to overfitting the domains on which they were RPT-trained, thereby amplifying in-domain improvements without corresponding generalization benefits. We present detailed results and analysis in Appendix C.
@@ 5 Related Work
RPT algorithms and frameworks. RPT, especially RLVR, has received significant attention in building powerful reasoning models, following the release of OpenAI o1 (OpenAI et al., 2024) and DeepSeek-R1 (DeepSeek-AI et al., 2025). RPT models have proven effective for a broad range of tasks with well-defined correctness: ranging from structured tasks such as math (Xiong et al., 2025; Cui et al., 2025) and coding (Wei et al., 2025; Shojaee et al., 2023; Le et al., 2022; Shen and Zhang, 2024), to unstructured tasks like search engine use (Jin et al., 2025) and open-ended question answering (Su et al., 2025a). However, all these models are trained and evaluated on tasks within a single domain or task type. Even works on general knowledge (Su et al., 2025a) remain confined to open-ended question answering tasks, without testing transfer across fundamentally different task types. In contrast, our work directly addresses this cross-domain generalization gap.
Limitations of RPT. Despite the recent successes of RPT in improving language model reasoning capabilities, the limitations of RPT in general have been widely studied. At the training phase, SFT with LLM reasoning traces, without RL, has been shown to be effective enough (Team, 2025b; Yue et al., 2025). At the inference phase, the quality of reasoning models depends crucially on their ability to scale under test-time compute constraints (Muennighoff et al., 2025; Zhao et al., 2025b). Moreover, the effectiveness of LLMs’ lengthy “thinking” processes has also been challenged (Ma et al., 2025; Ye et al., 2025). Building on these observations, our work aims to examine the limitations of RPT at a finer granularity, specifically its data generalizability across the training and inference phases.